A ground state hydrogen atom absorbs a photon of light having a wavelength of 92.05 nm. It then gives off a photon having a wavelength of 901.3 nm. What is the final state of the hydrogen atom?
1
Expert's answer
2016-11-05T11:31:07-0400
The process follows so-called Ridberg formula 1/l = R*(1/n1^2 - 1/n2^2):
1) 92.05 nm is a wavelength corresponded to n=1 to n=10 transferring of an electron (n2 = sqrt(Rl/Rl - 1) = 10);
2) 901.3 nm corresponds to transferring of an electron from 10th level to the 3th level (n1 = sqrt(100*Rl/(100 + Rl)) = 3)

Thus the final state is n=3 or the third quantum level.